OpenStudy (anonymous):

I Need a smarty ;) Heelp Pleasee. The equation of line GH is y = 4x − 3. Write an equation of a line perpendicular to line GH in slope-intercept form that contains point (2, 3).

OpenStudy (anonymous):

y = 4x − 3 y = 4x − 5 y = negative 1 over 4x − 3 y = negative 1 over 4x + 7 over 2

OpenStudy (anonymous):

its the last one right

OpenStudy (benlindquist):

uh yes

OpenStudy (anonymous):

thhhaaankkss